nts) Consider the fllowing two Security X and Y 1.2 Return 9.40% 1360% Risk-free assets 25% (1) Which security (X or Y) in the table hs risk? e has the least total risk? e least total risk? Which has the least che ecst third invested in Y? is the systematic risk for a portfolio with twor two-thirds of the funds invested in X and one- X, 45 (3) What is the portfolio beta and the portfolio expected return if you invest 35 percent in percent in Y, and 20 percent in the risk-free asset if expected market return is 8 20
Answer #1

Total risk = standard deviation, Systematic risk = beta

1) Security having the least standard deviation, i.e., Stock Y has the least Total risk. Security having the least beta, i.e., Stock X has the least systematic risk.

2) Portfolio systematic risk = 1.2 x 2/3 + 1.8 x 1/3 = 1.4

3) Portfolio beta = 1.2 x 0.35 + 1.8 x 0.45 + 0 x 0.20 = 1.23

Expected portfolio return = 9.40% x 0.35 + 13.60% x 0.45 + 1% x 0.20 = 9.61%